OpenStudy (anonymous):

Solve and show all work: |2x + 3| - 6 = 11

OpenStudy (anonymous):

You want to isolate the x. So add 6 to both sides.

OpenStudy (anonymous):

That leaves you with what?

OpenStudy (anonymous):

\[\left| x-a \right|=b,x-a=\pm b\]

OpenStudy (anonymous):

|2x+6| = 17 The number inside the absolute value sign can be positive or negative, so you have 2x+6 = 17 or 2x+6 = -17